Computing Double-Declining-Balance Depreciation mpany purchased equipment for $50,000. For each of the following sets of assumptions a depreciation schedule tall years) for this equipment assuming that DeFond uses the prepare Useful life a. Four years b. Five years c. Ten years Residual Value $8,000 $3,000 $1,000 Round answers to the nearest whole number. Use rounded answers for subsequent calculations. a. Year Book Value Depreciation Expense
media%2Fc1d%2Fc1ddf299-7ded-4ba9-807f-84
Answer #1

As per double declining depriciation method, rate of depriciation will be twice as much as it would have been under straight line method and then depriciation is to be charged on beginning book value of the period . Book value can be reduced till residual value only . Excess depriciation is then ignored.

a). Book value and depriciation expense will be as follows

Life of the asset = 4 years.

Depriciation rate as per SLM = (1/4)x100 = 25%.

Thus, as per double declining balance depriciation method , depriciation rate = 25%x2 = 50%.

Year book value depriciation exp
1 $25000 (50000-25000) 25000 (50000x50%)
2 $12500 (25000-12500) 12500 (25000x50%)
3 $8000

4500 (12500x50% restricted upto salvage value i.e 12500-8000 = $4500)

4 $8000 0

b). Useful life of asset = 5 years

Depriciation rate as as per SLM = (1/5)x100

= 20%

Depriciation as per double declining Balance depriciation method = 2x20% = 40%.

Years book value depriciation exp
1 $30000 (50000-20000) $20000 (50000x40%)
2 $18000 $12000 (30000x40%)
3 $10800 $7200 (18000x40%)
4 $6480 $4320 (10800x40%)
5 $3888 $2592 (6480x40%)

C) . Useful life = 10 years

Depriciation rate as per SLM = (1/10)x100 = 10%

Depriciation rate as per double declining method = 2x10% = 20%

Year Book value depriciation expense
1 $40000 $10000
2 $32000 $8000
3 $25600 $6400
4 $20480 $5120
5 $16384 $4096
6 $13107 $3277
7 $10486 $2621
8 $8389 $2097
9 $6711 $1678
10 $5369 $1342